OpenStudy (anonymous):

Is my answer correct? D: Alexander deposits $70 monthly into an account which pays 2.7% compounded semi-annually. How much remains in the account after 3 years?

OpenStudy (anonymous):

P = 70 i = 0.027 / 2 = 0.0135 n = 3(2) = 6 A = P(1 + i)^n = 70(1 + 0.0135)^6 = 70(1.0135)^6 = 75. 86

OpenStudy (anonymous):

Yep it's right! Doin' them now at school ;)